# Open-World Skill Discovery from Unsegmented Demonstrations

### v1

@@ 3.3 Implemetation Details
GROOT original 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 -
@@
STEVE-1 original 0.0% 1.0% 33.3% 33.3% 18.8% 65.6% 96.9% 18.8% 80.2% 57.3% 44.8% 46.9% -
@@
$\Delta$ 0 -- +21.9% +131.3% +127.8% +9.5% 0 +155.6% +5.2% +18.2% -2.3% +53.3% +52.1%
Table 2: Success rate of different policies on Minecraft skill benchmarks. For VPT (Baker et al., 2022), we report the results of the behavioral cloning version. For GROOT (Cai et al., 2023b) and STEVE-1 (Lifshitz et al., 2023), we report the results of original and our re-trained with SBD, respectively. A value with % indicates the average success rate, while a value without % indicates the average rewards. The seeds for the Minecraft environment are fixed for the corresponding task to make a fair comparison between different models. Details of the tasks are provided in Section B.1.
Policies for Atomic Skills
We employ both language-conditioned and video-conditioned policies to learn atomic skills from the segmented demonstration videos as shown in Stage III. The video-conditioned policy is built upon GROOT (Cai et al., 2023b), which utilizes self-supervised learning to model instructions and behaviors. In the original GROOT, a fixed-length sequence of 128 frames serves as the instruction, which is encoded into a goal embedding using a conditioned variational autoencoder (C-VAE) (Kingma and Welling, 2014; Sohn et al., 2015). The decoder then predicts actions based on the instruction and environmental observations in an auto-regressive manner. The language-conditioned policy is derived from STEVE-1 (Lifshitz et al., 2023), an instruction-tuned VPT model capable of following open-ended texts or 16-frame visual instructions. This model is trained by first adapting the pretrained VPT model to follow commands in MineCLIP’s (Fan et al., 2022) latent space, and then training a C-VAE model to predict latent codes from the text. The original GROOT and STEVE-1 policies divide the training trajectories into segments of 128 frames and uniformly sample between 15 and 200 frames, respectively. We re-train these models using our SBD method on the segmented trajectories and compare the results to the original models to demonstrate the effectiveness of our approach. We retain the minimum and maximum length settings. The algorithm for pruning the length of each trajectory is detailed in B.4. Most of the original hyperparameter settings are retained as specified in the original papers. A complete list of modified hyperparameters is provided in Section B.2.
@@
(a) OmniJARVIS (behavior cloning)
@@
Table 3: Success rate of two agents with their corresponding controllers trained on dataset segmented by our SBD method and the original sequential sampling (SS) method on groups of long programmatic tasks. All tasks are tested 30 times. In each group, the agent is required to obtain a certain type of items from scratch or given an iron pickaxe. For example, the diamond group includes diamond pickaxe, diamond sword, jukebox, etc.
Task [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Average
- 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 64.2
Info 33.0% 52.0% 100.0% 88.0% 97.0% 32.0% 23.7 65.0% 92.0% 47.0% 15.5 21.6 87.4
Loss 44.0% 54.0% 100.0% 94.0% 72.0% 46.0% 25.8 63.0% 95.0% 48.0% 17.9 22.7 91.8
Both 30.0% 54.0% 100.0% 88.0% 93.0% 80.0% 26.8 51.0% 90.0% 44.0% 19.7 25.4 93.3
Table 4: Ablation on the components within SBD. We report the evaluation results on Minecraft atomic skills from the sequential sampling (- in the table) and SBD with different components.
Long-term Instruction Following
@@
We use the hierarchical agents to model the long trajectories, which are widely used by (Driess et al., 2023; Wang et al., 2025, 2023b). We utilize in-context learning, as demonstrated by JARVIS-1 (Wang et al., 2024b), and imitation learning, as shown by OmniJARVIS (Wang et al., 2025), to evaluate the long-horizon instruction-following capabilities of hierarchical agents. JARVIS-1 is a hierarchical agent that uses the text-conditioned STEVE-1 as its policy and pretrained vision-language models as planners. We replace the text-conditioned policy with our re-trained STEVE-1 based on SBD. OmniJARVIS is a vision-language action agent built on FSQ-GROOT, which encodes instructions into discrete tokens rather than continuous embeddings. It is trained on behavior trajectories encoded into unified token sequences.
@@ 4 Experiments and Analysis
$\bullet$ The short segments obtained through SBD method, are they more consistent at the semantic level, and can a better atomic skill policy be trained on these segments?
$\bullet$ Does using short videos segmented with new methods as skills perform better on long-horizon tasks?
$\bullet$ Which component of SBD is the most important?
@@ 4.1 Experimental Setups
Environmental Setups and Datasets
We take the Minecraft environment (Guss et al., 2019; Zheng et al., 2023) as the evaluation simulator. Within the Minecraft environment, we use OpenAI’s contractor dataset 7.x (early game) (Baker et al., 2022) as trajectories. The dataset contains offline trajectories with 68M frames with a duration of approximately 1000 hours, with at least half of the data from the first 30 minutes of the game. Our method generates a segmented dataset of 130k sub-trajectories using bc-early-game-3x (Baker et al., 2022) as the pretrained unconditional model.
@@
For the atomic conditioned policies, we selected basic skills such as chop down trees and advanced skills like smelt items with furnace in Minecraft as evaluation benchmarks. We tested 12 different skill sets designed in MCU (Lin et al., 2023). Each task was tested over 100 times, except for sleep in bed and use bow, which were evaluated 10 times using the human Elo rating. The evaluation metrics are computed with the success rates and rewards. The hierarchical agents are evaluated on long-horizon programmatic tasks requiring the agents to start from an empty inventory in a new world until obtaining the final required items, such as obtain an iron pickaxe from scratch, which is usually a chain of atomic tasks. We split the long-horizon tasks into different groups, including wooden items, food, stones items, iron etc. For each group, we choose the tasks from JARVIS-1 benchmarks (Wang et al., 2024b) and evaluate each task over 10 times.
@@ 4.2 Main Results
In this section, we present the results of our method, SBD, applied to the two policies and two agents mentioned earlier. SBD achieves an average performance increase of 63.7% for GROOT and 52.1% for STEVE-1 compared to the original methods. Additionally, it enhances the performance of OmniJARVIS and JARVIS-1 by 11.3% and 20.8%.
@@
The results are shown in Table 2. Scores with % indicate success rates, while those without % represent rewards for the corresponding tasks. The controllers showed substantial improvements across most tasks, with an average performance enhancement of 63.7% for video-conditioned policy and 52.1% for language-conditioned policy.
@@
To further verify that the improvements in controllers enhance the ability of hierarchical agents, we evaluate OmniJARVIS and JARVIS-1 on programmatic tasks selected from the original papers (Wang et al., 2025, 2024b). As shown in Table 3, the agents have substantial improvements across most of the tasks, achieving an average performance enhancement of 11.3% for Omnijarvis and 20.8% for JARVIS-1.
4.3 Ablation Study
@@
Skill Videos
Figure 3: Video Segment Examples. Top: sleep in bed. Bottom: collect grass. Each segment is accompanied by five screenshots. The first and last screenshots represent the initial and final frames of the segment, respectively. The remaining three screenshots are manually selected to best illustrate the skill’s progression. More segments can be found in Appendix C.
As introduced in Sections 3.2.1 and 3.2.2, the environment information and the skill boundary detection loss are important components for accurate segmentation of long videos. In this section, we explore the effectiveness of every component. So we compare the full SBD with sequential sampling (w/o loss and info), SBD w/ info only, and SBD w/ loss only. To evaluate the effectiveness of these configurations, we conducted an ablation study on the performance of four methods under GROOT. The results are presented in Table 4. Our findings are as follows: (i) Using external information alone significantly improves performance, with an average increase of 36.1% over the original agent. (ii) Using loss alone results in a better performance (an average increase of 5.0%) compared to using external information alone, highlighting the effectiveness of the core component of our method. This demonstrates that our method achieves strong performance even without external information. (iii) Combining loss and external information yields the best overall performance. This combination outperforms the individual components in half of the tasks, suggesting that each component captures unique segmentation patterns that the other misses.
We observe performance declines when employ both components in tasks smelt food and collect seagrass. This may be due to both components partitioning the trajectory at the same intervals but not at identical steps, leading to redundant splits and incorrect trajectory fragments.
@@ 4.4 Visualizations
Fig. 2 shows the length distribution of sub-trajectories split by different methods with or without the two components of SBD. We observe that both loss-only and info-only methods follow similar distributions, where the log length and frequency of sub-trajectories exhibit a normal distribution. Since the info-only method is intrinsically semantically meaningful, this provides indirect evidence that the loss-only method also identifies a semantically meaningful segmentation pattern as the info-only method. Furthermore, when combining loss with external information, the distribution does not significantly differ from the loss-only method. This observation highlights that the predictive loss is the key driver in learning the segmentation pattern, with the external information serving a more supplementary role.
We sample one long unsegmented video segmented by our method for each skill in the datasets, such as sleep in bed and collect seagrass (see Fig. 3). More segments can be seen in Appendix C.
@@ 5 Related Works
Prior work on learning from unsegmented demonstrations has proposed various methods to segment trajectories into sub-trajectories. For instance, some methods (Shankar and Gupta, 2020; Tanneberg et al., 2021; Kipf et al., 2019) utilize a variational autoencoder (Kingma and Welling, 2014) model to generate pairs of skill types and durations, $(g_{i},t_{i})_{i=1}^{k}$, from a given trajectory, thereby enabling segmentation. TACO (Shiarlis et al., 2018) adopts a weakly supervised framework that identifies task boundaries, $(t_{1},\ldots t_{k})$ based on task sketches, $(g_{1},\ldots g_{k})$, by solving a sequence alignment problem to achieve appropriate segmentation. BUDS (Zhu et al., 2022) constructs hierarchical task structures of demonstration sequences using a bottom-up strategy, deriving temporal segments through agglomerative clustering of the actions. FAST (Pertsch et al., 2025) proposes an approach based on byte-pair encoding (Gage, 1994) algorithm for segmentation and tokenization of robot action trajectories via time-series compression.
@@
Option learning (Sutton et al., 1999) is a framework designed to learn temporal abstractions (i.e., skills or primitives) directly from the environment’s reward signal, primarily through online reinforcement learning. Recent approaches (Bacon et al., 2017; Bagaria and Konidaris, 2019; Klissarov and Precup, 2021) rely on policy gradient methods to learn the option policies. This framework requires reward signals, and DIAYN (Eysenbach et al., 2019) proposes a method to generate a reward function adaptively in the absence of reward signals, resulting in the unsupervised emergence of diverse skills.
@@
In neuroscience, episodic memory is the memory of everyday events consisting of short slices of experience (Conway, 2009). Event Segmentation Theory (EST) (Zacks et al., 2007) offers a theoretical perspective on how the neurocognitive system splits a long flow of memory into short events. According to EST, observers build event models of the current situation to generate predictions of future perceptual input. When errors in predictions arise, an event boundary is perceived, causing the event model to be reset and rebuilt. Our method employs a pretrained model to predict the agent’s action based on past observations and uses the predictive loss as an indicator to detect skill boundaries.
@@
Many agents have been developed to interact with Minecraft environments (Zhao et al., 2024a; Jiang et al., 2025). For short-horizon tasks, methods typically employ imitation or reinforcement learning, as seen in works like VPT (Baker et al., 2022), which annotated a large YouTube Minecraft video dataset with actions and trained the first foundation agent in the domain using behavior cloning, and its derivatives (Lifshitz et al., 2023; Cai et al., 2023b; Yuan et al., 2024; Cai et al., 2023a, 2024b). For long-horizon, programmatic tasks, large language models (LLMs) are used as planners combined with skill policies (Wang et al., 2023b; Zhou et al., 2024; Yuan et al., 2023; Qin et al., 2024; Wang et al., 2024b; Li et al., 2024), and some methods (Wang et al., 2024b; Zhu et al., 2023; Wang et al., 2024a) employ explicit memory mechanisms to further enhance the long-horizon capabilities of LLM agents. Additionally, recent advances (Wang et al., 2025; Zhao et al., 2024b) have explored using end-to-end visual-language models (VLMs) to directly follow human instructions. Finally, some research (Guo et al., 2024; Zhang et al., 2023) focuses on open-ended creative tasks such as building and decoration, which often cannot be directly defined by rule-based rewards.
6 Conclusions
In this paper, we propose a novel temporal video segmentation algorithm to address the challenges of open-world skill discovery from unsegmented demonstrations. To generate a segmented dataset of video clips with independent skills, our algorithm detects the potential skill boundaries based on the predictive loss of a pretrained action-prediction model. External boundary indicators can also be incorporated in our method. The method does not require any manual annotations and can be employed directly on massive internet gameplay videos.
7 Limitations and Future work
The current algorithm is relatively time-consuming, requiring approximately 1 minute of processing on an NVIDIA RTX 4090Ti for every 5-minute video, due to the need to predict actions and compute loss across the entire trajectory. As a result, we have not tested our method on larger datasets. Moreover, evaluating the method requires reprocessing the entire dataset and retraining the controllers, so we have not conducted an extensive ablation study on pretrained model selection or hyperparameter tuning. Additionally, we observe that the algorithm becomes unstable in action-intensive scenarios (e.g., combat situations), often generating many short segments. Future work could focus on improving the efficiency and stability of the algorithm, as well as exploring more effective evaluation methods. We also anticipate the potential application of our method to larger datasets and other open-world video games beyond Minecraft.

### v2

@@ 3.2.2 Boundary with External Information
GROOT
(video
conditioned)
original 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 -
@@
STEVE-1
(image & text
conditioned
original 0.0% 1.0% 33.3% 33.3% 18.8% 65.6% 96.9% 18.8% 80.2% 57.3% 44.8% 46.9% -
@@
$\Delta$ 0 - +21.9% +131.3% +127.8% +9.5% 0 +155.6% +5.2% +18.2% -2.3% +53.3% +52.1%
Table 2: Success rate of different policies on Minecraft skill benchmark. For VPT [3], we report the results of the behavioral cloning version. For GROOT [6] and STEVE-1 [27], we report the results of the original and our re-trained models with SBD, respectively. A value with % indicates the average success rate, while a value without % indicates the average rewards. See more details in Sec.˜D.1.
@@
(a) OmniJARVIS (behavioral cloning)
@@
Table 3: Success rate of two agents on long programmatic tasks. The goal-conditioned policies of the agents are trained on the dataset segmented by SBD and the original random splitting method. In each group, the agent is required to obtain a certain type of item from scratch or be given an iron pickaxe. For example, the diamond group includes diamond pickaxe, diamond sword, jukebox, etc.
3.3 Implemetation Details
We take the Minecraft environment [18, 49] as the evaluation simulator. It offers a highly complex (> $10^{10^{20}}$ states), diverse, and dynamic environment ideal for testing open-ended agents requiring OOD generalization [28]. See Appendix˜C for more information on this environment.
We now introduce the architectures and training/inference details of the policies and agents used in Sec.˜3.1.
Unconditional Policy and Segmentation Details
For the unconditional policy $\pi_{\text{unconditional}}$ in Stage I, we use the pre-trained vpt-3x models [3]. This is a foundational Transformer-XL [11] based model pretrained on large-scale YouTube data and finetuned on the early-game dataset using behavioral cloning. The hyperparameter GAP in Stage II is set to 18, considering the average trajectory length and semantics. As for the external information, we use event-based information such as mine_block:oak_log or use_item:torch available in the contractor dataset [3]. To identify potential skill transitions, only the final event in a series of identical consecutive events is marked as positive, as it signifies a possible skill transition. For instance, if the agent chops a tree repeatedly and then crafts a table, the video should be split at the moment the agent mines the last block of wood. Further details are provided in Sec.˜D.3.
Policies for Atomic Skills
We employ both video-conditioned and language-conditioned policies to learn atomic skills from the segmented demonstration videos, as shown in Stage III. The video-conditioned policy is built upon GROOT [6], which utilizes self-supervised learning to model instructions and behaviors. In the original GROOT, a fixed-length sequence of 128 frames serves as the instruction, which is encoded into a goal embedding using a conditioned variational autoencoder (C-VAE) [23, 33]. The decoder then predicts actions based on the instruction and environmental observations in an auto-regressive manner. The language-conditioned policy is derived from STEVE-1 [27], an instruction-tuned VPT model capable of following open-ended texts or 16-frame visual instructions. This model is trained by first adapting the pretrained VPT model to follow commands in MineCLIP’s [15] latent space, and then training a C-VAE model to predict latent codes from the text.
The original GROOT and STEVE-1 policies divide the training trajectories into segments of 128 frames and uniformly sample between 15 and 200 frames, respectively. We re-train these models using our SBD method on the segmented trajectories and compare the results to the original models to demonstrate the effectiveness of our approach. We retain the minimum and maximum length settings. The algorithm for pruning the length of each trajectory is detailed in D.4. Most of the original hyperparameter settings are retained as specified in the original papers. A complete list of modified hyperparameters is provided in Sec.˜D.2.
Long-term Instruction Following Agents
@@ 3.3 Implemetation Details
We use hierarchical agents to model long trajectories, which are widely used by [12, 40, 38]. We utilize in-context learning, as demonstrated by JARVIS-1 [39], and behavioral cloning, as shown by OmniJARVIS [40], to evaluate the long-horizon instruction-following capabilities of hierarchical agents. JARVIS-1 is a hierarchical agent that uses the text-conditioned STEVE-1 as its policy and pretrained vision-language models as planners. We replace the text-conditioned policy with our re-trained STEVE-1 based on SBD. OmniJARVIS is a vision-language action agent built on FSQ-GROOT, which encodes instructions into discrete tokens rather than continuous embeddings. It is trained on behavior trajectories encoded into unified token sequences.
@@ 4 Experiments and Analysis
Task
use
furnace
hunt
sheep
sleep
in bed
use
torch
use
boat
use
bow
collect
stone
collect
seagrass
collect
wood
find and
collect wood
collect
dirt
collect
grass
Average
- 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 63.3
good dist. 38.0% 65.0% 90.0% 89.0% 80.0% 10.0% 24.8 43.0% 96.0% 41.0% 15.5 19.0 79.9
Info 33.0% 52.0% 100.0% 88.0% 97.0% 32.0% 23.7 65.0% 92.0% 47.0% 15.5 21.6 85.8
Loss (GAP=17.5) 44.0% 49.0% 100.0% 89.0% 91.0% 32.0% 27.3 65.0% 96.0% 38.0% 17.8 22.7 88.4
Loss (GAP=18) 44.0% 54.0% 100.0% 94.0% 72.0% 46.0% 25.8 63.0% 95.0% 48.0% 17.9 22.7 90.2
Loss (GAP=18.5) 41.0% 51.0% 100.0% 91.0% 73.0% 44.0% 26.8 57.0% 91.0% 45.0% 18.1 25.3 88.5
Both (GAP=18) 30.0% 54.0% 100.0% 88.0% 93.0% 80.0% 26.8 51.0% 90.0% 44.0% 19.7 25.4 91.7
Table 4: Ablation studies. We report the evaluation results of GROOT trained on datasets segmented by random splitting with fixed length (-), random splitting following the length distribution of SBD (good dist.), and SBD with different components and GAP values. A value with % indicates the average success rate, while a value without % indicates the average rewards. To enable a fair comparison between success rate and rewards, the average score is computed by scaling the best score to 100 and all other scores proportionally for each task.
@@
$\bullet$ Are the short segments obtained through SBD more consistent semantically, and can a better goal-conditioned policy be trained on these segments?
$\bullet$ Do improvements in goal-conditioned policies enhance the ability of hierarchical agents?
$\bullet$ Which component of SBD is more important?
@@ 4.1 Experimental Setups
Training Dataset
We use OpenAI’s contractor dataset 7.x (early game) [3] as our training dataset. It contains Minecraft offline trajectories with 68M frames with a duration of approximately 1000 hours, with at least half of the data from the first 30 minutes of the game. Our method generates a segmented dataset of 130k sub-trajectories using bc-early-game-3x [3] as the pretrained unconditional model.
@@
For the goal-conditioned policies, we select basic skills such as collect wood and advanced skills like use furnace in Minecraft as the evaluation benchmark. We test 12 different atomic skills designed based on MCU [28]. All tasks are tested over 100 times, except for sleep in bed and use bow, which are evaluated 10 times using human rating because they cannot be automatically verified with in-game information. The evaluation metrics are success rates or rewards (for finer-grained assessment of easier tasks). See more details in Sec.˜D.1.
The hierarchical agents are evaluated on long-horizon programmatic tasks that require the agents to start from an empty inventory in a new world until obtaining the final required items, such as obtain an iron pickaxe from scratch, which is usually a chain of atomic tasks. We split the long-horizon tasks into different groups, including wooden items, food, stone items, iron etc. For each group, we choose the tasks from JARVIS-1 benchmarks [39] and evaluate each task over 30 times.
@@ 4.2 Main Results
In this section, we first present a human analysis of the quality of the SBD-segmented videos. Then we present the results of SBD applied to the two policies and two agents in Sec.˜3.3 on the benchmarks in Sec.˜4.1.
Human Analysis of Segmented Videos
We ask 10 people to analyze the SBD-segmented videos, each with 50 randomly sampled clips: On average, 54% exhibit clear independent semantics aligned with human preference, 34% are considered acceptable, and only 12% could be better segmented. This further supports our method’s effectiveness.
@@
As shown in Tab.˜2, the policies showed substantial improvements across most tasks, with an average relative performance enhancement of 63.7% for GROOT and 52.1% for STEVE-1.
@@
As shown in Tab.˜3, the agents have substantial improvements across most of the tasks, achieving an average relative performance enhancement of 11.3% for Omnijarvis and 20.8% for JARVIS-1.
4.3 Ablation Studies
@@
Figure 3: Video Segment Examples. Top: sleep in bed. Bottom: collect grass. Each segment is accompanied by five screenshots. The first and last screenshots represent the initial and final frames of the segment, respectively. The remaining three screenshots are manually selected to best illustrate the skill’s progression. More segments can be found in Appendix˜E.
SBD Components
As introduced in Secs.˜3.2.1 and 3.2.2, the environment information and the skill boundary detection loss are important components for accurate segmentation of long videos. In this section, we explore the effectiveness of each component. We compare the full SBD with random splitting (w/o loss and info), SBD w/ info only, and SBD w/ loss only. The results are presented in Tab.˜4. Our findings are as follows: (i) Using external information alone significantly improves performance, with a relative average increase of 35.5% over the original agent. (ii) Using loss alone results in a better performance (a relative average increase of 5.1%) compared to using external information alone, highlighting the effectiveness of the core component of our method. This demonstrates that our method achieves strong performance even without external information. (iii) Combining loss and external information yields the best overall performance. This suggests that each component captures unique segmentation patterns that the other misses.
We observe performance declines when employing both components in tasks use furnace and collect seagrass. This may be due to both components splitting the trajectory at the same intervals but not identical steps, leading to redundant trajectory fragments.
Hyperparameter sensitivity
GAP is an important hyperparameter of SBD. To verify the stability of our method, we also conduct an ablation study on different values of GAP. As shown in Tab.˜4, our method is not sensitive to it (only 2.0% relative average performance decline).
Improved Baseline
A simple idea to improve random splitting is to follow a similar length distribution as SBD. We add this improved baseline to show that our method also catches features other than length distribution. As shown in Tab.˜4, while a better length distribution can improve baseline performance (by 26.2% relatively), our method still outperforms it (by 14.8% relatively).
@@ 4.4 Visualizations
Fig.˜2 shows the length distribution of sub-trajectories split by different methods with or without the two components of SBD. We observe that both loss-only and info-only methods follow similar distributions, where the log length and frequency of sub-trajectories exhibit a normal distribution. Since the info-only method is intrinsically semantically meaningful, this provides indirect evidence that the loss-only method also identifies a semantically meaningful segmentation pattern as the info-only method. Furthermore, when combining loss with external information, the distribution does not significantly differ from the loss-only method. This observation highlights that the predictive loss is the key driver in learning the segmentation pattern, with the external information serving a more supplementary role.
Skill Videos
We sample one long unsegmented video segmented by our method for each skill in the datasets, such as sleep in bed and collect seagrass (see Fig.˜3). More segments can be seen in Appendix˜E.
@@ 5 Related Works
Prior work on learning from unsegmented demonstrations has proposed various methods to segment trajectories into sub-trajectories. For instance, some methods [31, 35, 24] utilize a variational autoencoder [23] model to generate pairs of skill types and boundaries, $(g_{i},t_{i})_{i=1}^{k}$, from a given trajectory. TACO [32] adopts a weakly supervised framework that identifies skill boundaries, $(t_{1},\ldots t_{k})$, based on skill sketches, $(g_{1},\ldots g_{k})$, by solving a sequence alignment problem. BUDS [52] constructs hierarchical task structures of demonstration sequences using a bottom-up strategy, deriving temporal segments through agglomerative clustering of the actions. FAST [29] proposes an approach based on the byte-pair encoding [16] algorithm for segmentation and tokenization of robot action trajectories.
@@
Option learning [34] is a framework to learn temporal abstractions (i.e., options or skills) directly from the environment’s reward signal, primarily through online reinforcement learning. Recent approaches [1, 2, 25] rely on policy gradient methods to learn the option policies. DIAYN [14] proposes a method to generate a reward function adaptively in the absence of reward signals, resulting in the unsupervised emergence of diverse skills.
@@
In neuroscience, episodic memory is the memory of everyday events consisting of short slices of experience [10]. Event Segmentation Theory (EST) [44] offers a theoretical perspective on how the neurocognitive system splits a long flow of memory into short events. According to EST, observers build event models of the current situation to generate predictions of future perceptual input. When errors in predictions arise, an event boundary is perceived, and the event model is reset and rebuilt. Our method employs a pretrained model to predict the agent’s action based on past observations and uses the predictive loss as an indicator to detect skill boundaries.
@@
Many agents have been developed to interact with Minecraft environments [47, 21]. For short-horizon tasks, methods typically employ behavioral cloning or reinforcement learning, as seen in works like VPT [3], which annotates a large Minecraft YouTube video dataset with actions and trains the first foundational agent in the domain, and its derivatives [27, 6, 43, 5, li2025jarvis]. For long-horizon, programmatic tasks, large language models (LLMs) are used as planners combined with skill policies [38, 50, 42, 30, 39, 26, ui-tars-15-seed], and some methods [39, 51, 36] employ explicit memory mechanisms to enhance the long-horizon capabilities of LLM agents. Additionally, recent advances [40, 48] have explored using end-to-end vision-language models (VLMs) to directly follow human instructions. Finally, some research [17, 45] focuses on open-ended creative tasks such as building and decoration.
6 Conclusion
In this paper, we propose a novel temporal video segmentation algorithm to address the challenges of open-world skill discovery from unsegmented demonstration videos. To generate a segmented dataset of video clips with independent skills, our algorithm detects the potential skill boundaries based on the predictive loss of a pretrained action-prediction model. It does not require any manual annotations and can be employed directly on massive internet gameplay videos.